The hydroxide ion concentration in an aqueous solution at 25°C is 8.9x10 The hydronium ion concentration is The pH of this solution is The pOH is M M. Rety Entme Group 9 more group attempts remaining Submit Answer
Answer #1

Kw (water) = 1.011 * 10-14

Kw = [H3O+] * [OH-] (product of Hydronium and Hydroxide ion concentrations.

Given [OH-] = 8.9 * 10-3M

Therefore,  [H3O+] = Kw / [OH-] = 1.011 * 10-14 /  8.9 * 10-3 = 0.1136 * 10-11 M = 1.136 * 10-12 M

Now pH = -log10[H3O+] = -log101.136 * 10-12 = 11.9

pOH = 14 - pH =14 - 11.9 = 2.1